Solve the system by substitution:2x-y=-5 and y=x+4
Answer by stanbon(75887)

Solve the system by substitution:2x-y=-5 and y=x+4
-----
2x -(x+4) = -5
2x -x - 4 = -5
-----
x = -1
--------
Cheers,
Stan H.
----------